Add & Subtract Mixed Numbers
(1 3/25)+ (8 2/15) but I have simplify?

Answer by JulietG(1812)

You'll need to first find a common denominator
25,50,75,100, etc
15,30,45,60,75,90,105, etc
You can see that the least common is 75, so we'll use that
1 3/25 is 1 9/75 (multiplied both parts of the fraction by 3)
8 2/15 is 8 10/75 (multiplied both parts of the fraction by 5)
Convert the mixed numbers into fractions
1 9/75 = 75/75 + 9/75 = 84/75
8 10/75 = 8*75/75 + 10/75 = 610/75
Add the like terms
84/75 + 610/75 = 694/75
Reduce
9 19/75